OpenStudy (ccastudentgrade10):

PLEASE HELP FAN AND MEDAL Arnie says that you can take the square root of each side of an equation. Therefore, he says that ( x + 2)2 + ( y – 5)2 = 64 and ( x + 2) + ( y – 5) = 8 are equivalent equations. Frankie says the equations are not equivalent. Who is right? Explain your answer. -I don't think you can take the square of both sides of the equation, but I don't understand how to find the answer??

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you can take the square root of both sides, but you would get \[\sqrt{(x+2)^2+(y-5)^2}=8\]

OpenStudy (misty1212):

and \[\sqrt{(x+2)^2+(y-5)^2}\]is not the same thing as \[(x+2)+(y-5)\] at all

OpenStudy (ccastudentgrade10):

okay, so Frankie is right? @misty1212 (sorry i only saw this now opensutdy didnt alert me!)